OpenStudy (anonymous):

solve this system using elimination x+y=7 x-y=3

OpenStudy (texaschic101):

x + y = 7 x - y = 3 ----------add 2x = 10 -- divide by 2 x = 5 now sub 5 in for x in either of the equations to find y x + y = 7 5 + y = 7 y = 7 - 5 y = 2 solution is (5,2)

OpenStudy (texaschic101):

any questions ?